### v1

@@ 1 Introduction
In today’s era, many media forms, including online news platforms, television channels, forums, blogs, and podcasts, serve as conduits of information to the public. These media continuously produce and disseminate news, presenting a complex challenge for researchers aiming to understand news impact [1]. Despite these challenges, multiple studies across various disciplines have employed techniques for analyzing news for diverse purposes. These include analyzing trends in news media [2], detecting bias in news articles [3], evaluating the effect of news on the public [4], developing tools to recognize fake news [5, 6] and more. For example, Vasiliki et al. [7] estimated countries’ peace index out from the Global Data on Events, Location, and Tone (GDELT) dataset. [^1] https://www.gdeltproject.org/ To perform that evaluation, they utilized news documents that were related to events by a machine-coded approach. This study is one example of many others who utilized event-based data to combat real-world issues [8, 9, 10]. As Liang Zhao stated ”Event analytics are important in domains as different as healthcare, business, cybersphere, politics, and entertainment, influencing almost every corner of our lives” [11]. Moreover, Zhao further claimed that ”The analysis of events has thus been attracting huge attention over the past few decades” [11].
@@
In their simplest form, event data convert natural language reports to a dataset where each entry has the form of Date, Source_Actor, Target_Actor, Event_Code [12]. Historically, many researchers conducted their studies using human-coded data, such as The Conflict and Peace Data Bank (COPDAB), which is a longitudinal computer-based library of daily international and domestic events or interactions [13] and WEIS Project dataset - which is a record of the flow of action and response between countries reflected in public facts [14]. However, these methods relied on costly human coding efforts that, after decades of research, came to an end [15]. As a result, event data started to be generated through content analysis - using pattern recognition and simple grammatical parsing methods made it possible to develop automated coding systems. An example of such a system is the Conflict and Mediation Event Observations (CAMEO), a widely used taxonomy [16]. The advancement toward automatic coding systems allowed researchers who are interested in employing event data to study international behavior and led to the development of computational methods for the analysis and prediction of various social, political, and economic studies [17, 18]. However, the automated machine-coded approaches are far from perfect as they are highly restrictive and still suffer from low accuracy, as suggested by Leetaru and Schrodt [15] - ”Small changes in the text or the parsing method could heavily affect the document’s tagged events.”
@@
To unlock the potential of news articles for these purposes, the development of meaningful news embeddings became a central challenge in the NLP research field, as suggested by Ma et al. [19] ”In order to facilitate the performance of downstream NLP tasks, it is rather essential to find an efficient way to represent news as continuous vectors.” Over the years multiple studies from different domains utilized news embedding in their work, some notable examples of such works are media bias prediction [3], fake news detection [20], and online news recommendations [21].
@@
Two main approaches to news embedding emerged: The first approach involves considering the complete information of articles. The prevailing methods in news embedding usually involve capturing semantic context with complex Deep Neural network architectures. For example, Baly et al. [3] used a BERT encoder to generate its news embeddings for its media bias classification task. The second approach relies on other features, such as topic modeling, NERs, and other semantical information, to approach the problem differently. News2vec [19] is an example of such an approach, in this work, news article embeddings were generated by incorporating relationship information between articles and events and using additional topological features in a graph comprised of extracted words and features from the article textual data and associated events.
@@
However, a major limitation of news embedding approaches is the ambiguous manner in which the embedding is constructed. News embeddings that provide good performance on several Natural Language Processing (NLP) tasks that do not necessarily succeed in association articles to events [22]. Moreover, efficiency plays a major part when constructing online news embeddings, which is usually a streaming task. Most of these works [3, 21, 23, 24] rely on textual network embedding methods and complex Deep Neural Networks (DNN) which are shown to be expensive both in memory and computation [25].
@@
In this study, we address these limitations by focusing on a specific facet and creating a vector space optimized for event-centric embeddings. To accomplish our goal, we relied on a combination of Named Entity Recognition (NER) techniques and topic modeling. This allows us to make accurate predictions while using fewer resources and providing clear results. To accomplish this, we introduce in this work a novel method for News Article Event-Based Embedding which receives as input themes, entities, and tagged events of a large set of articles, and the output is event-based embeddings on the given news article information.
@@
    Articles and Entities Processing - A large dataset of articles and entities is processed by executing multiple tasks, including removing redundancies, eliminating article duplication, applying source-based filtering, and entity parsing and removal.
@@
    Entities Embedding Generation - Given the processed dataset we trained separate word embedding models on current and historical articles to create informative embeddings for entities and themes. In this stage, each entity and theme receive a vector representation based on their co-occurrences with each other.
@@
    Articles Embedding Generation - given the entity embeddings, we deploy a Smooth Inverse Frequency (SIF) as a methodology to construct document-level vectors (see Section 2). Then, these embeddings traverse Siamese Neural Networks trained to minimize the distance between articles that share common events.
@@
To test and evaluate our study, we deployed our algorithm on the GDELT project, by collecting five years worth of online news article information from a predefined set of media sources across the United States - After processing the data contained more than 850,000 unique articles and more than one million events. We tested our hypothesis by measuring the proposed models’ ability to predict whether a pair of articles discuss common events in two different settings: daily test sets and monthly test sets. We measured both the Precision-Recall Area Under the Curve (PR AUC) metric and the Receiver Operating Characteristic Area Under the Curve (AUC). The results demonstrated a significant difference between our suggested semi-supervised method which relied on event-tagged historical data, and the baseline approach: On daily aggregated test datasets: our approach yields 0.572 PR-AUC, while the baseline yields 0.546 on average. On monthly aggregated test datasets: our best approach yields 0.369 PR-AUC, while the baseline yields 0.347 on average.
@@
  • We provide access to the research community for time-relevant embeddings of more than 320,00 distinct themes and entities. Time-relevant entity embeddings were generated depending on periodical corpora, that purposely overrepresent recently published news articles.
  • We suggest a new approach for evaluating the performance of news embedding methods, by pair-wise articles’ event attribution, which utilizes diverse real-world data of relationships between articles and events.
@@ 2 Related Work
This section contains the background necessary to understand the building blocks and motivation of the research. We discuss previous works in NLP Embedding, topic modeling, Named Entity Recognition, and studies on the GDELT project.
@@ 2.1.1 Word Embeddings Methods
In this work, we relied directly on the low-rank approximation of co-occurrences of themes and entities by using the GloVe model developed by Pennington et al. [26]. In their study, Pennington et al. proposed efficiently leveraging statistical information by training only on the nonzero elements in a word-word co-occurrence matrix rather than on the entire sparse matrix or individual context windows in a large corpus. In their results, GloVe provided better results than Word2Vec, and notably with higher efficiency, as they stated: ”the efficiency with which the count-based methods capture global statistics can be advantageous”  [26]. Therefore, GloVe can easily be adapted to the use case in this work since we did not utilize the news article’s full text but considered an efficient approach to use themes and entities while disregarding their position and textual context.
@@ 2.2 Topic Modeling
Understanding the core themes associated with a document collection is a fundamental task in today’s information era [35]. The purpose of topic models is to detect those ”themes,” which are the ideas and subjects that ”connect” a set of text documents. An example of a possible theme is ”police brutality.” These models are a class of unsupervised machine-learning techniques designed for this task. In this study, we define a topic model to be an unsupervised mathematical model that takes as input a set of documents $D$ and returns a set of topics $T$ that represent the content of $D$ accurately and coherently as suggested by Kherwa and Bansal [36].
A survey by Churchill and Singh [37] mentioned several studies that employed topic models for climate change literature [38], to understand the 2016 US presidential election using newspapers [39], additional works focused on developing new methods for topic modeling while evaluating their performance on news-based datasets such as 20 Newsgroups dataset [40, 41]. [^2] Link for the 20 Newsgroup dataset: http://qwone.com/~jason/20Newsgroups/
@@ 2.3 Named Entity Recognition
Named Entity Recognition (NER) aims to recognize mentions of rigid designators from text belonging to predefined semantic types such as person, location, organization, etc.” [42]. Named Entity Recognition (NER) is a versatile tool that enables the extraction of information from various sources. It is an essential component of Natural Language Processing (NLP) applications such as text comprehension [43], information retrieval [44], machine translation, and knowledge base construction [45].
The most commonly used typologies consist of a few high-level classes such as Person, Organisation, and Location [46]. There is a variety of open-source code designed for NER tasks, such as spaCy [47], NLTK [48], and flair [49]. A NER system aims to learn patterns from labeled examples, which can be used to classify new sequences of tokens.
@@
In our research, we relied on the NER algorithm output deployed by the GDELT project as input. They deployed three engines for recognizing persons, organizations, and locations by identifying an array of other kinds of proper names while attempting to compile a list of all amounts of each entity expressed in each article to offer numeric context to global events. Further description of this process can be found in the GDELT project codebook. [^3] http://data.gdeltproject.org/documentation/GDELT-Global_Knowledge_Graph_Codebook-V2.1.pdf
@@ 2.4 NLP News Embeddings
In 2018, Setty et al. presented Event2vec [22]. They used proposed network embedding techniques for the embedding representation of news events. As events involve different classes of nodes, such as named entities, temporal information, etc, general-purpose network embeddings are agnostic to event semantics. To address this problem, they proposed biased random walks tailored to capture news events’ neighborhoods in event networks.
@@
In 2019, Ma et al. presented News2vec [19], which represents the context of articles, events, and additional features in a graph. They focus on both the contextual relationship between news and potential connections as the network becomes more profound, with the embedded vectors containing semantic and labeled information, as well as latent connections between different news events.
@@
Since then, most of the works utilized news embeddings as part of a technique to achieve specific tasks [20, 21, 24], rather than suggesting a generic approach for news article embedding. For example, in 2022, Mehta et al. [20] work on fake detection included the following methodology. They embedded using graph neural network articles; then they formulated inference operators that augment the graph edges by revealing unobserved interactions between its elements, such as the similarity between documents’ contents and users’ engagement patterns.
@@
However, most of these work for news textual network embedding methods still rely on full-text and complex Deep Neural Networks, which are expensive both in memory and computation, as Xiang and Wang claimed in their event extraction study: ”How to design an efficient neural network architecture is the main challenging issue for event extraction based on deep learning” [54]. Furthermore, the works that handle the previous limitations usually fail to collect features that do not solely describe the contextual information at the document level as stated in [19] - ”The major limitation of embedding approaches discussed above is that they produce representations and features that solely describe the contextual information at the document level”. Another limitation is that in previous works in the field, the evaluation process depends on measuring several tasks that do not necessarily determine the success of associating articles with events. In the works of Ma et al. and Setty et al. [19, 22] that concentrated on the embeddings process of news, they chose for evaluation tasks of news recommendation, or stock prediction tasks, as well as the majority of works in the field which utilized news embeddings for a specific task, such as described in the following services [5, 11, 55, 23].
@@
There are numerous challenges to performing the task of associating events with articles, First, the positive rate is very low. In our case, on average the probability of two articles sharing a common event given they were published in the same month is less than 0.05% (see Table LABEL:tab:datasets_statistics). Second, An article can be associated with multiple events, creating a state of uncertainty in the detection process, and limiting the possible solutions. To alleviate this difficulty some previous works even chose to assume that a single article can be only related to one event [56]. Third, as we discussed earlier, robust methods that rely on full textual data, transformers, or graph neural networks often require high computational and memory resources [25, 57].
@@ 2.5 GDELT Project
The GDELT Project offers a comprehensive Event Database that contains records of more than 300 types of physical activities taking place around the world. These activities can range from riots and protests to peace appeals and diplomatic exchanges. The database provides the location of the activity, whether it’s a city or a mountaintop, and covers the entire planet, dating back to January 1, 1979. It is updated every 15 minutes, making it a reliable source of real-time information. Examples of some of the identified events include ”ENVIRONMENT AND NATURAL RESOURCES”, ”LABOR MARKETS” and ”CRISIS WELLBEING HEALTH”. [^6] https://blog.gdeltproject.org/gdelt-2-0-our-global-world-in-realtime/. GDELT also considered ’mentions’ - which are the unique pairings of the article and its discussed event. Multiple articles may discuss a shared event, and one article may discuss multiple events.
@@ 3.1.1 Entities and Themes Extraction
Given a corpus of textual articles $A$, we represent each article as a set of themes and entities. First, we suggest extracting those features from the textual corpus by deploying NER algorithms that extract locations, persons, dates, organizations, and the number of times each entity appeared in a given article. Additionally, we utilize topic models to extract topics and tones conveyed in the article.
@@ 3.1.2 Processing Articles Entities
We have implemented a naive method for performing entity matching. Namely, given the entity type, a set of normalizing operations was performed. We cast dates in a unified format. [^10] We cast dates to the following formats: Day/Month/Year (dd/mm/yyyy) or Day/Month (dd/mm). We cast persons and organizations to lower cases and locations to match their country and city codes. Next, we aggregated mappers for each entity type $<Entity,Count>$ during the pipeline. Lastly, when an entity that has not appeared yet is given as input, we check for the closest entity by measuring the hamming distance with a predefined threshold. Hamming distance is a popular measure of the distance or similarity between objects. The hamming distance is defined as the number of positions at which the corresponding symbols differ [59]. If the new entity does not have a match, it is inserted as a new entity to the relevant mapper.

### v2

@@ 1 Introduction
In today’s era, many media forms, including online news platforms, television channels, forums, blogs, and podcasts, serve as conduits of information to the public. These media continuously produce and disseminate news, presenting a complex challenge for researchers aiming to understand news impact [1]. Despite these challenges, multiple studies across various disciplines have employed techniques for analyzing news for diverse purposes. These include analyzing trends in news media [2], detecting bias in news articles [3], evaluating the effect of news on the public [4], developing tools to recognize fake news [5, 6], and more. For example, Vasiliki et al. [7] estimated countries’ peace index from the Global Data on Events, Location, and Tone (GDELT) dataset. [^1] https://www.gdeltproject.org/ To perform that evaluation, they utilized news documents that were related to events using a machine-coded approach. This study is one example of many others who have utilized event-based data to combat real-world issues [8, 9, 10]. As Liang Zhao stated ”Event analytics are important in domains as different as healthcare, business, cybersphere, politics, and entertainment, influencing almost every corner of our lives” [11]. Moreover, Zhao further claimed that ”The analysis of events has thus been attracting huge attention over the past few decades” [11].
@@
In its simplest form, event data converts natural language reports to a dataset where each entry has the form of Date, Source_Actor, Target_Actor, Event_Code [12]. Historically, many researchers conducted their studies using human-coded data, such as the Conflict and Peace Data Bank (COPDAB), which is a longitudinal computer-based library of daily international and domestic events or interactions [13], and the WEIS Project dataset, which is a record of the flow of action and response between countries as reflected in public facts [14]. However, these methods relied on costly human coding efforts that, after decades of research, came to an end [15]. As a result, event data started to be generated through content analysis using pattern recognition and simple grammatical parsing methods, making it possible to develop automated coding systems. An example of such a system is Conflict and Mediation Event Observations (CAMEO), a widely used taxonomy [16]. The advancement toward automatic coding systems allowed researchers who are interested in employing event data to study international behavior and led to the development of computational methods for the analysis and prediction of various social, political, and economic studies [17, 18]. However, the automated machine-coded approaches are far from perfect as they are highly restrictive and still suffer from low accuracy, as suggested by Leetaru and Schrodt [15] - ”Small changes in the text or the parsing method could heavily affect the document’s tagged events.”
@@
To unlock the potential of news articles for these purposes, the development of meaningful news embeddings became a central challenge in the NLP research field, as suggested by Ma et al. [19] ”In order to facilitate the performance of downstream NLP tasks, it is rather essential to find an efficient way to represent news as continuous vectors.” Over the years, multiple studies from different domains have utilized news embedding in their work, some notable examples of such works are media bias prediction [3], fake news detection [20], and online news recommendations [21].
@@
Two main approaches to news embedding emerged: The first approach involves considering the complete information of articles. The prevailing methods in news embedding usually involve capturing semantic context with Deep Neural network architectures. For example, Baly et al.  [3] used a BERT encoder to generate its news embeddings for its media bias classification task. The second approach relies on other features, such as topic modeling, NERs, and other semantic information, to approach the problem differently. News2vec [19] is an example of such an approach. In this work, news article embeddings were generated by incorporating relationship information between articles and events, and using additional topological features in a graph comprised of extracted words and features from the article textual data and associated events.
@@
However, a major limitation of news embedding approaches is the ambiguous manner in which the embedding is constructed. News embeddings that provide good performance on several Natural Language Processing (NLP) tasks do not necessarily succeed in the association of articles to events [22]. Moreover, efficiency plays a major part when constructing online news embeddings, since it is considered a streaming task. Most of these works [3, 21, 23, 24] rely on textual network embedding methods and complex Deep Neural Networks (DNN), which are shown to be expensive both in memory and computation [25].
@@
In this study, we address these limitations by focusing on a specific facet and creating a vector space optimized for event-centric embeddings. To accomplish our goal, we relied on a combination of Named Entity Recognition (NER) techniques and topic modeling. This allows us to make accurate predictions while using fewer resources and providing clear results. To accomplish this, we introduce in this work a novel method for News Article Event-Based Embedding that receives as input themes, entities, and tagged events of a large set of articles, and the output is event-based embeddings on the given news article information.
@@
    Articles and Entities Processing: A large dataset of articles and entities is processed by executing multiple tasks, including removing redundancies, eliminating article duplication, applying source-based filtering, and entity parsing and removal.
@@
    Entities Embedding Generation: Given the processed dataset, we trained separate word embedding models on current and historical articles to create informative embeddings for entities and themes. In this stage, each entity and theme receives a vector representation based on their co-occurrences with each other.
@@
    Articles Embedding Generation: Given the entity embeddings, we deploy a Smooth Inverse Frequency (SIF) as a methodology to construct document-level vectors (see Section 2). Then, these embeddings traverse Siamese Neural Networks trained to minimize the distance between articles that share common events.
@@
To test and evaluate our study, we deployed our algorithm on the GDELT project, by collecting five years worth of online news article information from a predefined set of media sources across the United States. After processing, the data contained more than 850,000 unique articles and more than one million events. We tested our hypothesis by measuring the proposed models’ ability to predict whether a pair of articles discuss common events in two different settings: daily test sets and monthly test sets. We measured both the Precision-Recall Area Under the Curve (PR AUC) metric and the Receiver Operating Characteristic Area Under the Curve (AUC). The results demonstrated a significant difference between our suggested semi-supervised method, which relied on event-tagged historical data, and the baseline approach. On daily aggregated test datasets, our approach yields 0.572 PR-AUC, while the baseline yields 0.546 on average. On a monthly aggregated test datasets, our best approach yields 0.369 PR-AUC, while the baseline yields 0.347 on average.
@@
  • We provide access to the research community for time-relevant embeddings of more than 320,00 distinct themes and entities. Time-relevant entity embeddings were generated, depending on periodical corpora that purposely overrepresent recently published news articles.
  • We suggest a new approach for evaluating the performance of news embedding methods: pair-wise articles’ event attribution, which utilizes diverse real-world data on relationships between articles and events.
@@ 2 Related Work
This section contains the background necessary to understand the building blocks and motivation of the research. We discuss previous works in NLP embedding, topic modeling, named entity recognition (NER), and studies on the GDELT project.
@@ 2.1.1 Word Embeddings Methods
In this work, we utilized the GloVe model developed by Pennington et al. [26], relying directly on the low-rank approximation of co-occurrences of themes and entities. Pennington et al. proposed an efficient method of leveraging statistical information by training only on the nonzero elements in a word-word co-occurrence matrix, instead of the entire sparse matrix or individual context windows in a large corpus. Their results showed that GloVe outperformed Word2Vec and was more efficient, stating: ”the efficiency with which the count-based methods capture global statistics can be advantageous”  [26]. Since we did not use the full text of news articles, GloVe is well-suited for this work’s use case. Instead, we adopted an efficient approach that focuses on themes and entities while ignoring their position and textual context.
@@ 2.2 Topic Modeling
Understanding the core themes associated with a document collection is a fundamental task in today’s information era [35]. The purpose of topic models is to detect those ”themes,” which are the ideas and subjects that ”connect” a set of text documents. An example of a possible theme is ”police brutality.” These models are a class of unsupervised machine-learning techniques designed for this task. In this study, we define a topic model as an unsupervised mathematical model that takes as input a set of documents $D$ and returns a set of topics $T$ that represent the content of $D$ accurately and coherently as suggested by Kherwa and Bansal [36].
Churchill and Singh’s survey [37] cited several studies that used topic models to analyze climate change literature [38] and to understand the 2016 US presidential election through newspapers [39]. Additional works focused on developing new methods for topic modeling and evaluating their performance on news-based datasets, such as the 20 Newsgroups dataset [40, 41]. [^2] Link for the 20 Newsgroup dataset: http://qwone.com/~jason/20Newsgroups/
@@ 2.3 Named Entity Recognition
Named Entity Recognition (NER) is a versatile tool that enables the extraction of information from various sources. It is an essential component of Natural Language Processing (NLP) applications such as text comprehension [42], information retrieval [43], machine translation, and knowledge base construction [44].
The most commonly used typologies consist of a few high-level classes such as person, organization, and location [45]. There is a variety of open-source code designed for NER tasks, such as spaCy [46], NLTK [47], and flair [48]. A NER system aims to learn patterns from labeled examples, which can be used to classify new sequences of tokens.
@@
In our research, we relied on the NER algorithm output deployed by the GDELT project as input. They deployed three engines for recognizing emotions, dates, persons, organizations, and locations by identifying an array of other kinds of proper names while attempting to compile a list of all amounts of each entity expressed in each article to offer numeric context to global events. Further description of this process can be found in the GDELT project codebook. [^3] http://data.gdeltproject.org/documentation/GDELT-Global_Knowledge_Graph_Codebook-V2.1.pdf
@@ 2.4 NLP News Embeddings
In 2018, Setty et al. presented Event2vec [22]. They used proposed network embedding techniques for the embedding representation of news events. As events involve different classes of nodes, such as named entities, temporal information, etc. General-purpose network embeddings are agnostic to event semantics. To address this problem, they proposed biased random walks tailored to capture news events’ neighborhoods in event networks.
@@
In 2019, Ma et al. presented News2vec [19], which represents the context of articles, events, and additional features in a graph. They focus on both the contextual relationship between news and potential connections as the network becomes more profound, with the embedded vectors containing semantic and labeled information as well as latent connections between different news events.
@@
Since then, most of the works have utilized news embeddings as part of a technique to achieve specific tasks [20, 21, 24], rather than suggesting a generic approach for news article embedding. For example, in 2022, Mehta et al. [20] work on fake detection included the following methodology: They embedded using graph neural network articles; then they formulated inference operators that augment the graph edges by revealing unobserved interactions between its elements, such as the similarity between documents’ contents and users’ engagement patterns.
@@
However, most of these work for news textual network embedding methods still rely on full-text and complex DNN, which are expensive both in memory and computation, as Xiang and Wang claimed in their event extraction study: ”How to design an efficient neural network architecture is the main challenging issue for event extraction based on deep learning”  [53]. Furthermore, the works that handle the previous limitations usually fail to collect features that do not solely describe the contextual information at the document level, as stated in [19]: ”The major limitation of embedding approaches discussed above is that they produce representations and features that solely describe the contextual information at the document level”. Another limitation is that in previous works in the field, the evaluation process depends on measuring several tasks that do not necessarily determine the success of associating articles with events. In the works of Ma et al. and Setty et al. [19, 22] that concentrated on the embedding process of news, they chose for evaluation tasks of news recommendation, or stock prediction tasks, as well as the majority of works in the field that utilized news embeddings for a specific task, such as described in the following services [5, 11, 54, 23].
@@
There are numerous challenges to performing the task of associating events with articles. First, the positive rate is very low. In our case, on average, the probability of two articles sharing a common event given they were published in the same month is less than 0.05% (see Table LABEL:tab:datasets_statistics). Second, an article can be associated with multiple events, creating a state of uncertainty in the detection process and limiting the possible solutions. To alleviate this difficulty, some previous works even chose to assume that a single article could only be related to one event [55]. Third, as we discussed earlier, robust methods that rely on full textual data, transformers, or graph neural networks often require high computational and memory resources [25, 56].
@@ 2.5 GDELT Project
The GDELT Project offers a comprehensive event database that contains records of more than 300 types of physical activities taking place around the world. These activities can range from riots and protests to peace appeals and diplomatic exchanges. The database provides the location of the activity, whether it’s a city or a mountaintop, and covers the entire planet, dating back to January 1, 1979. It is updated every 15 minutes, making it a reliable source of real-time information. Examples of some identified events include ”ENVIRONMENT AND NATURAL RESOURCES”, ”LABOR MARKETS” and ”CRISIS WELLBEING HEALTH”. [^6] https://blog.gdeltproject.org/gdelt-2-0-our-global-world-in-realtime/. GDELT also considers ”mentions,” which are the unique pairings of an article and the event it discusses. Multiple articles can discuss the same event, and a single article can cover multiple events.
@@ 3.1.1 Entities and Themes Extraction
Given a corpus of textual articles $A$, we represent each article as a set of themes and entities. First, we suggest extracting those features from the textual corpus by deploying NER algorithms that extract locations, persons, dates, organizations, and the number of times each entity appears in a given article. Additionally, we utilize topic models to extract the topics and tones conveyed in the article.
@@ 3.1.2 Processing Articles Entities
We have implemented a naive method for performing entity matching. Namely, given the entity type, a set of normalizing operations was performed. We cast dates in a unified format. [^10] We cast dates to the following formats: Day/Month/Year (dd/mm/yyyy) or Day/Month (dd/mm). We cast persons and organizations into lower cases and locations to match their country and city codes. Next, we aggregated mappers for each entity type $<Entity,Count>$ during the pipeline. Lastly, when an entity that has not appeared yet is given as input, we check for the closest entity by measuring the hamming distance with a predefined threshold. Hamming distance is a popular measure of the distance or similarity between objects. The hamming distance is defined as the number of positions at which the corresponding symbols differ [58]. If the new entity does not have a match, it is inserted as a new entity to the relevant mapper.